∑ r=1 10 (n +1) =?? Why is it equal to 10(n+1) and not only (n+1)?? Please explain??

r=110 (n +1) =??
Why is it equal to 10(n+1) and not only (n+1)??
Please explain??

Grade:

1 Answers

Susmita
5 years ago
Hey i hope you know the formula 
\sum_{1}^{n}1 =n
Because \sum_{1}^{n}1means adding one n times.
In the given problem
 
\sum_{1}^{10}(n+1)
= \sum_{1}^{10}n +\sum_{1}^{10}1
= n\sum_{1}^{10}1 +\sum_{1}^{10}1
=n×10+10=10(n+1)
Hope this helps.please approve if it does.